# Geometry and Optimization of Shallow Polynomial Networks

### v1

@@ Appendix A Proofs and Additional Results
We now prove that any two minimal factorizations of $S$ are path connected in $\tau_{r}^{-1}(S)$ if ${\rm rk}(S)<r$. We first observe that any two minimal factorzations $(\alpha,W)$, $(\alpha^{\prime},W^{\prime})$ with $\alpha=\alpha^{\prime}$ are connected. Indeed, if $W^{+}\in\mathbb{R}^{\alpha_{+}\times n}$ and $W^{\prime+}\in\mathbb{R}^{\alpha_{+}\times n}$ are submatrices of $W$ and $W^{\prime}$ corresponding to elements of $\alpha$ that are $1$ (the same argument applies to $W^{-}\in\mathbb{R}^{\alpha_{-}\times n}$ and $W^{\prime-}\in\mathbb{R}^{\alpha_{-}\times n}$ corresponding to to elements of $\alpha$ that are $-1$), then we have that $Row(W^{+})=Row(W^{\prime+})$, so there exist $T\in GL(\alpha_{+})$ such that $W^{\prime+}=TW^{+}$. If $\det(T)>0$, then we can find a path $GL(\alpha_{+})$ from $Id_{\alpha_{+}}$ to $T$ that transforms $W^{\prime+}$ to $W^{+}$ without affecting $S$. If $\det(T)<0$, then we use a zero row of $W$ to change the sign of one row vector $w$ of $W^{+}$. More precisely, if $j$ is such that $\alpha_{j}=0,w_{j}=0$, then we consider a path that changes the sign of $w_{i}$ while keeping $\alpha_{i}(t)w_{i}(t)w_{i}(t)^{T}+\alpha_{j}(t)w_{j}(t)w_{j}(t)^{T}$ constant:
$\alpha_{i}(t)=\begin{cases}\sqrt{1-3t}\cdot\alpha_{i}&t\in[0,1/3]\\ 0&t\in[1/3,2/3]\\ (3t-2)\cdot\alpha_{i}&t\in[2/3,1]\\ \end{cases},$ $w_{i}(t)=\begin{cases}\sqrt{1-3t}\cdot w_{i}&t\in[0,1/3]\\ -\sqrt{3t-1}\cdot w_{i}&t\in[1/3,2/3]\\ -w_{i}\cdot\alpha_{i}&t\in[2/3,1]\\ \end{cases}$ (35)
$\alpha_{j}(t)=\begin{cases}\sqrt{3t}\cdot\alpha_{i}&t\in[0,1/3]\\ \alpha_{i}&t\in[1/3,2/3]\\ \sqrt{3-3t}\cdot\alpha_{i}&t\in[2/3,1]\\ \end{cases},$ $w_{j}(t)=\begin{cases}\sqrt{3t}\cdot w_{i}&t\in[0,1/3]\\ w_{i}&t\in[1/3,2/3]\\ \sqrt{3-3t}\cdot\alpha_{i}&t\in[2/3,1].\\ \end{cases}$
@@
If instead we assume $\alpha\neq\alpha^{\prime}$, then we observe the existence of a path in $\tau_{r}^{-1}(S)$ that changes $\alpha$ into $\alpha^{\prime}$. Indeed, there exists a path very similar to (35) such that $\alpha_{j}(0)=0$, $w_{j}(0)=0$, $\alpha_{j}(1)=\alpha_{i}$, $w_{j}(1)=w_{i}$ and $\alpha_{i}(1)=0$, $w_{i}(1)=0$ (it is enough to consider the subpath of (35) corresponding to $t\in[0,1/3]$). In other words, this path exchanges a zero row and a non-zero row. By composing similar paths, we can also exchange a row where the corresponding element of $\alpha$ is $1$ and one where the corresponding element of $\alpha$ is $-1$. Using these exchanges between rows, we can reduce the number of indices where $\alpha$ and $\alpha^{\prime}$ differ until $\alpha=\alpha^{\prime}$ holds. Finally, from the previous analysis, we deduce that when ${\rm rk}(S)<r$ the fiber $\tau_{r}^{-1}(S)$ is connected. This concludes the proof of the theorem. ∎
@@
We remark that the fiber has many more connected components over a matrices with mixed signature than over semi-definite ones. The number of connected components can be viewed as a local measure of complexity of the landscape.
@@
$\rho_{r}:\mathbb{R}^{r}\times V_{r}(\mathbb{R}^{n})\rightarrow{\rm Sym}^{2}(\mathbb{R}^{n}),\quad(\alpha,U)\mapsto U^{T}diag(\alpha)U.$
@@
$d\rho_{r}(\alpha,U)=\sum_{i=1}^{r}\dot{\alpha}_{i}(u_{i}\otimes u_{i})+\sum_{i=1}^{r}\alpha_{i}(u_{i}\otimes\dot{u}_{i}+\dot{u}_{i}\otimes u_{i}),$ (36)
[…]
@@
If $\Sigma$ corresponds to the Frobenius norm, then $\mathcal{N}_{X,\Sigma}=\{(S,A)\colon S\in X,\,S\cdot A=0\}$. Our assumptions on $S$ and $A$ imply that $S=\sum_{i=1}^{r}\sigma_{i}^{\prime}(u_{i}^{\prime}\otimes u_{i}^{\prime})$ and $A=\sum_{j=1}^{n-r}\sigma_{j}^{\prime\prime}(u_{j}^{\prime\prime}\otimes u_{j}^{\prime\prime})$ each have distinct nonzero eigenvalues. It follows now from Proposition 39 that $(\sigma^{\prime},\sigma^{\prime\prime},u^{\prime},u^{\prime\prime})\mapsto\left(\sum_{i=1}^{r}\sigma_{i}^{\prime}(u_{i}^{\prime}\otimes u_{i}^{\prime}),\sum_{j=1}^{n-r}\gamma\sigma_{j}^{\prime\prime}(u_{j}^{\prime\prime}\otimes u_{j}^{\prime\prime})\right)$ is a smooth local parameterization of $\mathcal{N}_{X,\Sigma}$ in a neighborhood of $(S,\gamma A)$. Since $u^{\prime},u^{\prime\prime}$ are orthogonal, another application of Proposition 39 implies that the nullity of the differential of the endpoint map $(S,\gamma A)\mapsto S+\gamma A$ is given by the number of pairs of repeated eigenvalues of $T_{\gamma}=S+\gamma A$.
If $\Sigma$ corresponds to the Gaussian norm, then $\mathcal{N}_{X,\Sigma}=\{(S,A)\colon S\in X,\,S\cdot(2A+{\rm tr}(A){\rm Id})=0\}$. Similar to the previous case, a smooth local parameterization of $\mathcal{N}_{X,\Sigma}$ in a neighborhood of $(S,\gamma A)$ is given by $(\sigma^{\prime},\sigma^{\prime\prime},u^{\prime},u^{\prime\prime})\mapsto\left(\sum_{i=1}^{r}\sigma_{i}^{\prime}(u_{i}^{\prime}\otimes u_{i}^{\prime}),-\sum_{i=1}^{r}\gamma c(u_{i}^{\prime}\otimes u_{i}^{\prime})+\sum_{j=1}^{n-r}\gamma\sigma_{j}^{\prime\prime}(u_{j}^{\prime\prime}\otimes u_{j}^{\prime\prime})\right)$ where $c=\frac{1}{r+2}\sum_{j=1}^{n-r}\sigma_{j}^{\prime\prime}$ and $u^{\prime}$ and $u^{\prime\prime}$ are orthogonal (see the proof of Theorem 35 for a derivation). It follows again from Proposition 39 the nullity of the differential of the endpoint map is the number of pairs of repeated eigenvalues of $T_{\gamma}=S+\gamma A$. Here we reduce ourselves to the previous setting by applying the invertible linear map $(\sigma^{\prime},\sigma^{\prime\prime})\mapsto(\sigma^{\prime}-\gamma c,\sigma^{\prime\prime})$ to obtain the eigenvalues of $T_{\gamma}$. ∎
@@
Automatically (39) holds at coordinates $i\in[n]$ where $v_{i}=0$. Thus we restrict (39) to the support of $v$, that is, =$\mathcal{I}=\{i:v_{i}\neq 0\}$. Assuming $\mu_{2}\neq 0$, (39) implies $v_{\mathcal{I}},v^{\ast 3}_{\mathcal{I}},t_{\mathcal{I}}\ast v_{\mathcal{I}}$ are linearly dependent, where subscripts indicate subvectors. Therefore for all $i,j,k\in\mathcal{I}$,
@@
$\mathcal{T}=\{t^{\prime}\in\mathbb{R}^{n}:\textup{ all coordinates of }t^{\prime}\textup{ are distinct and positive and }1\leq\frac{\max_{i\in[n]}t^{\prime}_{i}}{\min_{i\in[n]}t^{\prime}_{i}}\leq 2\},$ (51)
@@
It follows $\beta+\gamma t_{i}$ has sign $\varepsilon$ for all $i\in\mathcal{I}$. By (49) and (43), $\varepsilon=1$ and $v_{i}^{2}=\beta+\gamma t_{i}$ for $i\in\mathcal{I}$. Thus, there exist $\tau_{i}\in\{1,-1,0\}$ for $i\in[n]$ with $\tau_{i}\neq 0$ if and only if $i\in\mathcal{I}$ such that $v_{i}=\tau_{i}\sqrt{\beta+\gamma t_{i}}\in\mathbb{R}\setminus\{0\}$ where $\beta,\gamma$ are as in (48), and $S=\varepsilon vv^{\top}\in\mathbb{R}^{n\times n}$ has entries $\tau_{i}\tau_{j}\sqrt{(\beta+\gamma t_{i})(\beta+\gamma t_{j})}$. Conversely, such $S$ indeed satisfies the criticality condition (38). So depending on the choice of $\tau$, there is a critical point with support specified by $\mathcal{I}$. Choices $\tau$ and $-\tau$ give the same $S$ and otherwise the matrices $S$ are distinct. It follows there are $2^{|\mathcal{I}|-1}$ many critical points with support determined by $\mathcal{I}$.
The other case is $|\mathcal{I}|=1$. Then, $v$ is a scalar multiple of $e_{i}$, where =$\mathcal{I}=\{i\}$. Returning back to (40), it is easy to directly compute $S=\frac{\mu_{2}^{2}\langle t,1\rangle-\mu_{2}^{2}t_{i}+\mu_{4}t_{i}}{\mu_{4}}e_{i}e_{i}^{\top}$. This is real, nonzero and satisfies (38). So again, there are $2^{|\mathcal{I}|-1}$ many critical points in this case.
@@
Here, the second equality is by the binomial theorem. Finally, note that (51) has positive Lesbesgue-measure in $\mathbb{R}^{n}$. The proof of Theorem 36 is complete. ∎

### v2

@@ Appendix B Proofs and Additional Results
We now prove that any two minimal factorizations of $S$ are path connected in $\tau_{r}^{-1}(S)$ if ${\rm rk}(S)<r$. We first observe that any two minimal factorzations $(\alpha,W)$, $(\alpha^{\prime},W^{\prime})$ with $\alpha=\alpha^{\prime}$ are connected. Indeed, if $W^{+}\in\mathbb{R}^{\alpha_{+}\times n}$ and $W^{\prime+}\in\mathbb{R}^{\alpha_{+}\times n}$ are submatrices of $W$ and $W^{\prime}$ corresponding to elements of $\alpha$ that are $1$ (the same argument applies to $W^{-}\in\mathbb{R}^{\alpha_{-}\times n}$ and $W^{\prime-}\in\mathbb{R}^{\alpha_{-}\times n}$ corresponding to to elements of $\alpha$ that are 1-1), then we have that $Row(W^{+})=Row(W^{\prime+})$, so there exist $T\in GL(\alpha_{+})$ such that $W^{\prime+}=TW^{+}$. If $\det(T)>0$, then we can find a path in $GL(\alpha_{+})$ from $Id_{\alpha_{+}}$ to $T$ that transforms $W^{\prime+}$ to $W^{+}$ without affecting $S$. If $\det(T)<0$, then we use a zero row of $W$ to change the sign of one row vector $w$ of $W^{+}$. More precisely, if $j$ is such that $\alpha_{j}=0,w_{j}=0$, then we consider a path that changes the sign of $w_{i}$ while keeping $\alpha_{i}(t)w_{i}(t)w_{i}(t)^{T}+\alpha_{j}(t)w_{j}(t)w_{j}(t)^{T}$ constant:
[…]
@@
If instead we assume $\alpha\neq\alpha^{\prime}$, then we observe the existence of a path in $\tau_{r}^{-1}(S)$ that changes $\alpha$ into $\alpha^{\prime}$. Indeed, there exists a path very similar to (35) such that $\alpha_{j}(0)=0$, $w_{j}(0)=0$, $\alpha_{j}(1)=\alpha_{i}$, $w_{j}(1)=w_{i}$ and $\alpha_{i}(1)=0$, $w_{i}(1)=0$ (it is enough to consider the subpath of (35) corresponding to $t\in[0,1/3]$). In other words, this path exchanges a zero row and a non-zero row. By composing similar paths, we can also exchange a row where the corresponding element of $\alpha$ is 11 and one where the corresponding element of $\alpha$ is 1-1. Using these exchanges between rows, we can reduce the number of indices where $\alpha$ and $\alpha^{\prime}$ differ until $\alpha=\alpha^{\prime}$ holds. Finally, from the previous analysis, we deduce that when ${\rm rk}(S)<r$ the fiber $\tau_{r}^{-1}(S)$ is connected. This concludes the proof of the theorem. ∎
@@
We remark that the fiber has many more connected components over matrices with mixed signature than over semi-definite ones. The number of connected components can be viewed as a local measure of complexity of the landscape.
@@
$\rho_{r}:\mathbb{R}^{r}\times V_{r}(\mathbb{R}^{n})\rightarrow{\rm Sym}^{2}(\mathbb{R}^{n}),\quad(\alpha,U)\mapsto U^{T}\operatorname{diag}(\alpha)U.$
@@
$d\rho_{r}(\alpha,U)=\sum_{i=1}^{r}\dot{\alpha}_{i}(u_{i}\otimes u_{i})+\sum_{i=1}^{r}\alpha_{i}(u_{i}\otimes\dot{u}_{i}+\dot{u}_{i}\otimes u_{i}).$ (36)
From the condition $UU^{T}={\rm Id}_{r}$, the tangent space of $V_{r}(\mathbb{R}^{n})$ is characterized by $\dot{u}_{i}^{\top}u_{i}=0$ and $\dot{u}_{i}^{\top}u_{j}+\dot{u}_{j}^{\top}u_{i}=0$ for all $i\neq j$ in $\{1,\ldots,r\}$. Assuming without loss of generality that $U=[{\rm I}_{r}\,\,0]$, the image of (36) is an $n\times n$ symmetric matrix with $\dot{\alpha}_{i}$ in the first $r$ entries of the diagonal and $\alpha_{i}(u_{i}\otimes\dot{u}_{i})_{ij}+\alpha_{j}(\dot{u}_{j}\otimes u_{j})_{ij}=\alpha_{i}(\dot{u}_{i})_{j}+\alpha_{j}(\dot{u}_{j})_{i}=(\alpha_{i}-\alpha_{j})(\dot{u}_{i})_{j}$ in the $(i,j)$-th entry with $i\neq j$ in $\{1,\ldots,r\}$ (since $u_{i},u_{j}$ are rows of $U=[{\rm I}_{r},0]$ and $\dot{u}_{i}^{\top}u_{j}+\dot{u}_{j}^{\top}u_{i}=(\dot{u}_{i})_{j}+(\dot{u}_{j})_{i}=0$). From this description, the claim follows easily. ∎
@@
If $\Sigma$ corresponds to the Frobenius norm, then ${N_{X,\Sigma}}=\{(S,A)\colon S\in X,\,S\cdot A=0\}$. Our assumptions on $S$ and $A$ imply that $S=\sum_{i=1}^{r}\sigma_{i}^{\prime}(u_{i}^{\prime}\otimes u_{i}^{\prime})$ and $A=\sum_{j=1}^{n-r}\sigma_{j}^{\prime\prime}(u_{j}^{\prime\prime}\otimes u_{j}^{\prime\prime})$ each have distinct nonzero eigenvalues. It follows now from Proposition 40 that $(\sigma^{\prime},\sigma^{\prime\prime},u^{\prime},u^{\prime\prime})\mapsto\left(\sum_{i=1}^{r}\sigma_{i}^{\prime}(u_{i}^{\prime}\otimes u_{i}^{\prime}),\sum_{j=1}^{n-r}\gamma\sigma_{j}^{\prime\prime}(u_{j}^{\prime\prime}\otimes u_{j}^{\prime\prime})\right)$ is a smooth local parameterization of $N_{X,\Sigma}$ in a neighborhood of $(S,\gamma A)$. Since $u^{\prime},u^{\prime\prime}$ are orthogonal, another application of Proposition 40 implies that the nullity of the differential of the endpoint map $(S,\gamma A)\mapsto S+\gamma A$ is given by the number of pairs of repeated eigenvalues of $T_{\gamma}=S+\gamma A$.
If $\Sigma$ corresponds to the Gaussian norm, then $N_{X,\Sigma}=\{(S,A)\colon S\in X,\,S\cdot(2A+{\rm tr}(A){\rm Id})=0\}$. Similar to the previous case, a smooth local parameterization of $N_{X,\Sigma}$ in a neighborhood of $(S,\gamma A)$ is given by $(\sigma^{\prime},\sigma^{\prime\prime},u^{\prime},u^{\prime\prime})\mapsto\left(\sum_{i=1}^{r}\sigma_{i}^{\prime}(u_{i}^{\prime}\otimes u_{i}^{\prime}),-\sum_{i=1}^{r}\gamma c(u_{i}^{\prime}\otimes u_{i}^{\prime})+\sum_{j=1}^{n-r}\gamma\sigma_{j}^{\prime\prime}(u_{j}^{\prime\prime}\otimes u_{j}^{\prime\prime})\right)$ where $c=\frac{1}{r+2}\sum_{j=1}^{n-r}\sigma_{j}^{\prime\prime}$ and $u^{\prime}$ and $u^{\prime\prime}$ are orthogonal (see the proof of Theorem 36 for a derivation). It follows again from Proposition 40 the nullity of the differential of the endpoint map is the number of pairs of repeated eigenvalues of $T_{\gamma}=S+\gamma A$. Here we reduce ourselves to the previous setting by applying the invertible linear map $(\sigma^{\prime},\sigma^{\prime\prime})\mapsto(\sigma^{\prime}-\gamma c,\sigma^{\prime\prime})$ to obtain the eigenvalues of $T_{\gamma}$. ∎
@@
Automatically (39) holds at coordinates $i\in[n]$ where $v_{i}=0$. Thus we restrict (39) to the support of $v$, that is, $\mathcal{I}=\{i:v_{i}\neq 0\}$. Assuming $\mu_{2}\neq 0$, (39) implies $v_{\mathcal{I}},v^{\ast 3}_{\mathcal{I}},t_{\mathcal{I}}\ast v_{\mathcal{I}}$ are linearly dependent, where subscripts indicate subvectors. Therefore for all $i,j,k\in\mathcal{I}$,
@@
[…]
@@
It follows $\beta+\gamma t_{i}$ has sign $\varepsilon$ for all $i\in\mathcal{I}$. By (49) and (43), $\varepsilon=1$ and $v_{i}^{2}=\beta+\gamma t_{i}$ for $i\in\mathcal{I}$. Thus, there exist $\tau_{i}\in\{1,-1,0\}$ for $i\in[n]$ with $\tau_{i}\neq 0$ if and only if $i\in\mathcal{I}$ such that $v_{i}=\tau_{i}\sqrt{\beta+\gamma t_{i}}\in\mathbb{R}\setminus\{0\}$ where $\beta,\gamma$ are as in (48), and $S=\varepsilon vv^{\top}\in\mathbb{R}^{n\times n}$ has entries $\tau_{i}\tau_{j}\sqrt{(\beta+\gamma t_{i})(\beta+\gamma t_{j})}$. Conversely, such $S$ indeed satisfies the criticality condition (38). So depending on the choice of $\tau$, there is a critical point with support specified by $\mathcal{I}$. Choices $\tau$ and $\tau$ give the same $S$ and otherwise the matrices $S$ are distinct. It follows there are $2^{|\mathcal{I}|-1}$ many critical points with support determined by $\mathcal{I}$.
The other case is $|\mathcal{I}|=1$. Then, $v$ is a scalar multiple of $e_{i}$, where $\mathcal{I}=\{i\}$. Returning back to (40), it is easy to directly compute $S=\frac{\mu_{2}^{2}\langle t,1\rangle-\mu_{2}^{2}t_{i}+\mu_{4}t_{i}}{\mu_{4}}e_{i}e_{i}^{\top}$. This is real, nonzero and satisfies (38). So again, there are $2^{|\mathcal{I}|-1}$ many critical points in this case.
@@
Here, the second equality is by the binomial theorem. Finally, note that (51) has positive Lebesgue-measure in $\mathbb{R}^{n}$. The proof of Theorem 37 is complete. ∎
Appendix C Full polynomial from Example 29
$+\mu_{4}^{12}t_{00}^{3}t_{11}^{3}$